the ratio between sue and her brother sam is 3:7 if sam is 145 cm tall how tall is sue..

Answer by timvanswearingen(106)

The answer can be found by setting up a proportion and using cross products.
3/7 = x/145
Using cross products,
3*145 = 7*x
435=7x
Divide both sides by 7,
435/7 = 7x/7
62 1/7 = x
Sue is 62 1/7 cm tall.